# PointViG: A Lightweight GNN-based Model for Efficient Point Cloud Analysis

### v1

@@ 4.8.1 Visualization of t-SNE feature clustering
In Fig. 7 and Fig. 8, t-Distributed Stochastic Neighbor Embedding (t-SNE) is applied for both dimensionality reduction and cluster analysis on the encoder features and logits of the ModelNet40 test set samples. In Fig. 7, a significant portion of the clusters is clearly distinguishable from others, while a small central cluster lacks distinct boundaries, leading to some overlap. The introduction of logits, obtained by transforming encoder features through an MLP in the classification head, markedly clarifies the boundaries between clusters. In Fig. 8, apart from a limited number of outliers with abnormal distributions, points corresponding to the same category exhibit close proximity within their respective clusters. Notably, clusters associated with distinct categories are visibly distant from one another, emphasizing a high level of discriminability.
Figure 7: Visualisation of encoder features. We employed t-SNE to visualize the encoder features of the samples within the test set of ModelNet40.
Figure 8: Visualization of output logits. We utilize t-SNE to illustrate the logits of the samples within the ModelNet40 test set.

### v2

@@ 4.8.1 Visualization of t-SNE Feature Clustering
To further evaluate the feature representation capabilities of the proposed PointViG model, we employed t-SNE (t-Distributed Stochastic Neighbor Embedding) for the downscaling and clustering analysis of encoder features. This technique was applied to the ModelNet40 test set, enabling a comparative visual analysis between PointViG and the widely-used DGCNN model [9], which also adopts a GNN-based architecture. The clustering results are presented in Fig. 7(a) for DGCNN [9] and Fig. 7(b) for PointViG.
Figure 7: t-SNE visualization of encoded features for (a) DGCNN [9] and (b) PointViG on the ModelNet40 test set.
The t-SNE visualizations of DGCNN’s features reveal a pattern where the majority of clusters are not clearly separated, indicating that DGCNN [9] struggles to maintain distinct and independent feature representations. This suggests a limited capacity in the DGCNN [9] model to differentiate between similar point cloud structures within the dataset. The lack of well-defined boundaries highlights potential challenges in achieving robust classification performance.
In contrast, the feature space of PointViG, as visualized through t-SNE, demonstrates a substantial improvement in cluster separability. The clusters are independently distributed with minimal overlap, and most are tightly grouped, with clear boundaries between them. This pattern reflects PointViG’s enhanced ability to capture diverse geometric features from the point cloud data, leading to a more discriminative and robust feature representation. Such results suggest that the PointViG model is capable of better generalization and classification performance compared to DGCNN [9].